OpenStudy (pedro):

Write an expression for the phrase “the sum of ten and the product of a number and four.” List the key words and which operations they indicate.

OpenStudy (anonymous):

let eh no. be x product of no. and four = x*4 = 4x so the sum of ten and the product of a number and four = 10 + 4x or 4x + 10

OpenStudy (anonymous):

10+4x